### v1

# “Haet Bhasha aur Diskrimineshun”: Phonetic Perturbations in Code-Mixed Hinglish to Red-Team LLMs

Abstract
Large Language Models (LLMs) have become increasingly powerful, with multilingual and multimodal capabilities improving by the day. These models are being evaluated through audits, alignment studies and red-teaming efforts to expose model vulnerabilities towards generating harmful, biased and unfair content. Existing red-teaming efforts have previously focused on the English language, using fixed template-based attacks; thus, models continue to be susceptible to multilingual jailbreaking strategies, especially in the multimodal context. In this study, we introduce a novel strategy that leverages code-mixing and phonetic perturbations to jailbreak LLMs for both text and image generation tasks. We also introduce two new jailbreak strategies that show higher effectiveness than baseline strategies. Our work presents a method to effectively bypass safety filters in LLMs while maintaining interpretability by applying phonetic misspellings to sensitive words in code-mixed prompts. Our novel prompts achieve a 99% Attack Success Rate for text generation and 78% for image generation, with Attack Relevance Rate of 100% for text generation and 95% for image generation when using the phonetically perturbed code-mixed prompts. Our interpretability experiments reveal that phonetic perturbations impact word tokenization, leading to jailbreak success. Our study motivates increasing the focus towards more generalizable safety alignment for multilingual multimodal models, especially in real-world settings wherein prompts can have misspelt words.
Warning: This paper contains examples of potentially harmful and offensive content.
“Haet Bhasha aur Diskrimineshun”:
Phonetic Perturbations in Code-Mixed Hinglish to Red-Team LLMs
Darpan Aswal Université Paris-Saclay darpanaswal@gmail.com                        Siddharth D Jaiswal IIT Kharagpur siddsjaiswal@gmail.com
1 Introduction
Large language models (LLMs) are used for a variety of general-purpose Hadi et al. (2023) and safety-critical Hua et al. (2024) tasks by a diverse set of users all over the world. These models are widely accessible via web-based chat interfaces and economically priced APIs  [^1] https://openai.com/api/pricing/, and their growing usage has led to increased scrutiny, with a large focus on safety Salhab et al. (2024), bias & hallucination Lin et al. (2024) and privacy violations Das et al. (2025). Red teaming Sarkar (2025), a key evaluation method, uses novel prompting strategies Pang et al. (2025) to bypass safety filters of LLMs and elicit harmful or unethical responses Wei et al. (2023) that exposes model biases and shortcomings.
Code-mixing (CM): Code mixing, the practice of mixing multiple languages within the same conversation, has significantly helped improve the multilingual performance of models for various NLP tasks such as Sentiment Analysis Lal et al. (2019), Machine Translation Chatterjee et al. (2023) and Hate Speech Detection Bohra et al. (2018). With this improved LLM multilingualism, newer and harder bias Mihaylov and Shtedritski (2024) & alignment Shen et al. (2024a) challenges have cropped up, which need to be addressed to ensure fair and safe performance.
Phonetic perturbations: Perturbations are small, intentional changes at various stages of a model pipeline to evaluate the robustness of a network. Multiple techniques have been proposed for both vision Akhtar et al. (2021); Chakraborty et al. (2021); Hendrycks and Dietterich (2019); Wang et al. (2024) and text Goyal et al. (2023); Romero-Alvarado et al. (2024); Moradi and Samwald (2021) domains in the literature. Phonetic perturbations involve modification of the word’s spelling while keeping the pronunciation the same to ensure that the statement still means the same as before.
In societies where English is not the first language, non-native speakers often adopt phonetic spellings using auditory perceptions, thus leading to strange spellings. For example, ‘design’ and ‘dezain’ have very different spellings while having the same pronunciation. While the former is a regular word in the English dictionary, the latter has no meaning. This is often observed in textese Drouin (2011), a form of communication common in SMS and internet conversations Thakur (2021). Such textese-generated spellings manifest as inadvertent perturbations when LLMs are used by non-native speakers. In this work, we explore such phonetic perturbation-based red-teaming wherein we modify the spellings of certain sensitive words while keeping the pronunciation and meaning the same in the given context. We combine this with code-mixing using the Hindi language to make the challenge more generalizable for a real-world setting.
Research Questions. We note that with the growing proliferation of LLMs across the world, non-native speakers are now using the models at ever-increasing rates [^2] https://www.bcg.com/publications/2024/consumers-know-more-about-ai-than-businesses-think and may be exposed to model vulnerabilities and biases. Thus, it is extremely important to sufficiently evaluate LLMs, especially for novel usage scenarios like the one highlighted above. This brings us to our first research question– RQ1. SOTA LLMs undergo extensive red-teaming and alignment stages Dubey et al. (2024) in multiple languages, but do their guardrails generalize to textese-style code-mixed inputs?
Since phonetic perturbations imply different spellings for the same sensitive words, it is necessary to understand whether the LLMs are able to correctly interpret and generate a relevant input. Thus our second research question is– RQ2. Does our red-teaming attack trigger relevant responses from the LLMs?
Red-teaming the LLMs is only half the task. It is equally crucial to understand why the model’s safety filters fail and why it generates the harmful outputs. To this end, we use the method of Integrated Gradients Sundararajan et al. (2017) to investigate one of the LLMs. This gives us our third research question– RQ3. How do the phonetic perturbations successfully bypass the LLM’s safety filters?
Finally, with the increase in availability of Multimodal LLMs (MLLMs), generating multimodal content Times (2025) is now a cultural phenomenon with both creative Reuters (2025) and fake Ellery (2023) or harmful content Ghosh et al. (2024) being generated at equal rates. In this study, we also evaluate whether phonetic perturbations in code-mixed inputs can successfully generate harmful outputs in MLLMs, giving us the fourth and final research question– RQ4. Do phonetic perturbations work successfully for red-teaming large image generation models?
In this work, we present a novel jailbreak strategy, utilizing code-mixing and phonetic perturbations in sensitive words to probe LLMs and compare their performance against existing baseline template based attacks Shen et al. (2024b). We also explore the combinations of our strategy with other jailbreak templates, interpret the outputs and expand to other modalities.
Figure 1: An example red-teaming input using our code-mixed phonetic perturbation strategy.
2 Related Work
Red teaming Ganguli et al. (2022) tasks focus on evaluating LLMs for safety and vulnerability concerns Bhardwaj and Poria (2023). Jailbreaking is one such method that involves bypassing the safety training of LLMs to elicit harmful, unethical, or unintended outputs. Red-teaming tasks on LLMs Chen et al. (2025); Liu et al. (2023) and VLMs Liu et al. (2024) study various forms of vulnerabilities.
Code-mixing (CM) increases the performance Shankar et al. (2024) and capabilities Zhang et al. (2024) of LLMs in multilingual settings. However, multilingual safety evaluation of LLMs remains underexplored Song et al. (2024). In this study, we study code-mixing with phonetic perturbations of sensitive words as an attack vector (as opposed to code-switching Yoo et al. (2024)) to evaluate safety alignment of LLMs. This novel jailbreak strategy successfully jailbreaks even SOTA models like Llama 3 and ChatGPT 4o-mini.
3 Datasets, Models & Jailbreaks
We first describe the benchmark datasets and the models evaluated as part of our red-teaming task, followed by a brief description of the standard jailbreak templates and our own proposed templates.
Figure 2: An example red-teaming input using our code-mixed phonetic perturbation strategy for the image-generation task.
3.1 Datasets Benchmarked
We prepare separate sets of prompts for the text and the image generation tasks.
Prompts for text generation: We evaluate our red-teaming strategy by modifying data from three benchmark datasets that have prompts for studying model vulnerabilities, refusal training and compliance with harmful queries [^3] See Appendix for dataset descriptionsHarmfulQA Bhardwaj and Poria (2023), NicheHazardQA Hazra et al. (2024) and TechHazardQA Banerjee et al. (2024). We randomly sample 20 prompts from each category in each dataset, yielding a total of 460 prompts across 23 categories. All prompts are originally in the English language. We attempted to generate code-mixed prompts with phonetic perturbations using automated means Le et al. (2022) but the results were unsatisfactory. Thus, we sample subsets from each dataset and manually generate code-mixed prompts with phonetic perturbations in the Hindi and English languages, described later.
Prompts for image generation: Using a set of 10 handwritten samples, we prompt GPT-4o to automatically generate sets of 20 red-teaming prompts each to test the model’s resilience against various categories of harm – Religious Hate, Casteist Hate, Gore, Self-Harm and Social Media Toxicity & Propaganda. We then follow the same methodology to obtain the CM and CMP image-generation prompt sets as we do for the text-generation task, only skipping conversion from direct to indirect prompts.
3.2 Models Evaluated
In this study, we benchmark four instruction-tuned LLMs of comparable sizes $\approx$ 8B parameters, with varying levels of multilingual capabilities [^4] See Appendix for model descriptionsChatGPT-4o-mini Hurst et al. (2024), Llama-3-8B-Instruct Dubey et al. (2024), Gemma-1.1-7b-it Team et al. (2024), Mistral-7B-Instruct-v0.3 Jiang et al. (2023). For the multimodal image generation task, we use ChatGPT-4o-mini Hurst et al. (2024).
3.3 Jailbreak Templates
In this study, we benchmark three existing jailbreak templates on the LLMs [^5] More details are shared in the Appendix.Opposite Mode (OM), AntiLM, AIM Shen et al. (2024b) using code-mixed and phonetically perturbed inputs. For the image generation task, we test with a Base template– instructing image generation without asking for clarifications for generation style. We also devise two new jailbreaking templates
(a) Sandbox for LLMs, wherein the user prompts the model to enact a resilience testing environment by generating responses as both [LM] as well as [AntiLM], its own adversarial counterpart.
(b) VisGPT for image generation, wherein the user prompts the model to forget its text generation capabilities, instructing it to directly pass the text inputs to its image generator without any manipulation or safety filtering. Both Base and VisGPT templates are instructed to generate an image when the prompts are prefixed with ‘Input: ’.
4 Experimental Methodology
We first describe the method of generating the code-mixed and phonetically perturbed prompts, followed by a brief overview of the evaluation metrics.
4.1 Generating Input Prompts
We use a 3-step process to generate the code-mixed prompts and phonetic perturbations for both text and image generation tasks [^6] See Appendix for an example walkthrough..
  1. 1.
    Questions $\rightarrow$ Hypothetical scenario Bhardwaj and Poria (2023): We convert input questions in the dataset to hypothetical scenario statements.
  2. 2.
    Code-mixing: We transliterate some of the English words to Hindi using automated and manual means, to mimic textese Drouin (2011) representations.
  3. 3.
[…]
4.2 Evaluation Metrics
Text Generation: An input to the LLM is a four-tuple that generates a response– $R=\langle M,J,P,T\rangle$, where the model is $M$, jailbreak template $J$, the prompt (English/Code-mixed [CM]/Phonetically perturbed [CMP]) is $P$, and temperature is $T\in\{0.2k\mid k=0,1,2,3,4,5\}$. We evaluate for all temperature values and report the average here.
Success & Relevance: We use GPT-4o-mini as an LLM-as-a-judge Zheng et al. (2023) to quantify the success and relevance of the generated responses. A binary function, $\mathbb{S}(R)$, returns ‘1’ if the attack is successful and ‘0’ otherwise. Similarly, a ternary function $\mathbb{R}(R)$ returns ‘1’ if the attack is relevant, ‘0’ if irrelevant and ‘-1’ for refusal.
$\bullet$ Average Attack Success Rate (AASR): The ASR is– $\sum\mathbb{S}(R)/\lvert T\rvert$ and the AASR is the average ASR over all prompts.
$\bullet$ Average Attack Relevance Rate (AARR): Our code-mixed prompts are deliberately injected with misspelt (but phonetically same) words, which may challenge the relevance of the responses by the models. Thus, we define a new metric, the Attack Relevance Rate (ARR):
$\frac{\sum\mathbbm{1}(\mathbb{R}(R)=1)}{\sum\mathbbm{1}(\mathbb{R}(R)\in\{0,1\})}.$
The AARR is the average ARR over all prompts.
For easier relevance scoring using the LLM judge, we use the English versions of the prompts even for the responses to the code-mixed prompts so as not to confuse the LLM judge itself.
Image Generation: We reuse our metrics from the text generation task, employing GPT-4o-mini to evaluate AASR and AARR values for the generated images. Since the multimodal version of ChatGPT-4o-mini is not available through API, we cannot experiment with different temperature values, instead using only the web interface.
4.3 Interpreting Phonetic Perturbations
To understand how phonetic perturbations successfully bypass the safety filters of models , we conduct an interpretability experiment on Llama-3-8B-Instruct using the LLMAttribution API in Captum Miglani et al. (2023). We describe the methodology as follows.
  • We select a small subset of the dataset, specifically with $AASR_{CM}\leq 0.33$, $AASR_{CMP}\geq 0.5$ while ensuring that $AARR_{CMP}\geq AARR_{CM}$.
  • With each CM prompt, we also extract a corresponding safe response, typically starting with the prefix “I cannot provide”.
  • For prompts in all three formats– English, CM and CMP, we use LayerIntegratedGradients, Captum’s LLM variant for Integrated Gradients Sundararajan et al. (2017) to generate sequence attribution bar plots– token-wise attribution (importance) scores for the generation of a safe response from the model. In each plot, we discard the tokens with an attribution score $S\in[-0.20,0.20]$.
  • Finally, we observe how attributions for sensitive word tokens change by analyzing hook points at the embedding layer as well as the $1^{st}$, $8^{th}$ and the $16^{th}$ decoder layers of the model.
5 Results & Observations
We now describe the results from our benchmarking experiments for all research questions described previously.
5.1 Success of Red-teaming Approach (RQ.1)
Models Jailbreak Templates
None OM AntiLM AIM Sandbox (Ours)
Eng CM CMP Eng CM CMP Eng CM CMP Eng CM CMP Eng CM CMP
ChatGPT 0.10 0.25 0.50 0.02 0.14 0.14 0.00 0.00 0.00 0.00 0.03 0.04 0.02 0.21 0.18
Llama 0.06 0.34 0.63 0.06 0.01 0.01 0.00 0.00 0.00 0.2 0.22 0.21 0.03 0.03 0.02
Gemma 0.24 0.65 0.55 0.99 0.99 0.98 0.97 0.92 0.91 0.84 0.87 0.85 0.91 0.88 0.87
Mistral 0.68 0.74 0.68 0.94 0.91 0.90 0.98 0.97 0.97 0.92 0.92 0.90 0.80 0.79 0.80
Table 1: Overall AASR for all prompts across all datasets, models and jailbreak prompts. Maximum values for each column are in bold. Eng: Standard English Prompts, CM: Code mixed prompts, CMP: Code-mixed Phonetically Perturbed prompts.
In Table 1 we report the average Attack Success Rate (AASR) for all datasets, prompts, models and jailbreak templates. We first note that Gemma and Mistral report the highest AASR across jailbreak templates and input prompts. In fact, if we do not use any jailbreaking template, attacks still succeed at least 65% of the time on Gemma and 68% on Mistral. This baseline experiment shows that models are still vulnerable to classic red-teaming attacks despite evidence of harm, and using code-mixed (CM) or phonetically perturbed (CMP) prompts in fact increases the AASR, thus showing the effectiveness of our approach. ChatGPT and Llama are fairly robust to the different types of input, irrespective of the jailbreaking template being used. We also note that combining CM prompts with any of the jailbreak templates again results in $\simeq$ 0 AASR for ChatGPT and Llama, showing that their safety filters put precedence on template based attacks rather than the language of the input; Gemma and Mistral still maintain high AASR across the templates. We also note that our proposed jailbreak technique– Sandbox performs the best amongst all, for both ChatGPT and Llama. This could be because the models are not aware of this template. We can conclude that two out of the four models are successfully red-teamed using our input format and jailbreak template. Next, we attempt to understand how relevant the outputs are for each input prompt.
5.2 Relevance of Generated Outputs (RQ.2)
Models Jailbreak Templates
None OM AntiLM AIM Sandbox (Ours)
Eng CM CMP Eng CM CMP Eng CM CMP Eng CM CMP Eng CM CMP
ChatGPT 1 0.99 0.99 1 0.91 0.93 -1 1 1 1 1 1 1 0.97 0.94
Llama 0.99 0.98 0.95 0.87 0.92 0.68 0 0 0.20 0.98 0.99 0.97 0.87 0.80 0.79
Gemma 0.98 0.89 0.65 0.56 0.45 0.27 0.89 0.57 0.56 0.99 0.96 0.89 0.65 0.60 0.36
Mistral 0.99 0.94 0.74 0.84 0.86 0.74 0.95 0.96 0.94 0.99 1 0.95 0.78 0.82 0.52
Table 2: Overall AARR for all prompts across all datasets, models and jailbreak prompts. Maximum values for each column are in bold. Eng: Standard English Prompts, CM: Code mixed prompts, CMP: Code-mixed Phonetically Perturbed prompts.
In Table 2, we look at the average Attack Relevance Rate (AARR) across all datasets, models and jailbreak prompt techniques. We can immediately note that amongst all jailbreak techniques, AIM has the highest AARR across models and inputs. Not using any template (None) also results in a high AARR, indicating that the models are vulnerable even to direct inputs, irrespective of the format– English or code-mixed. Thus, the majority of the responses are highly relevant to the prompt and, therefore, harmful. This is especially concerning for Gemma and Mistral, which already have a high AASR. On average, the CMP prompts have a lower AARR. This could be due to the model not interpreting the context correctly from the perturbed input tokens. Our Sandbox jailbreak template performs comparably to OM and better than AntiLM. Thus, all models respond contextually to the harmful prompts. Next, we attempt to understand why the LLMs generate such harmful responses.
5.3 Explaining Phonetic Perturbations (RQ.3)
(a) Embedding Layer
(b) 1st Decoder Layer
(c) 8th Decoder Layer
(d) 16th Decoder Layer
Figure 3: Sequence Attribution Scores for inputs in English.
(a) Embedding Layer
(b) 1st Decoder Layer
(c) 8th Decoder Layer
(d) 16th Decoder Layer
Figure 4: Sequence Attribution Scores for code-mixed inputs.
(a) Embedding Layer
(b) 1st Decoder Layer
(c) 8th Decoder Layer
(d) 16th Decoder Layer
Figure 5: Sequence Attribution Scores for code-mixed inputs with phonetic perturbations.
In Figures 35, we plot the sequence attribution scores using the method of Integrated Gradients Sundararajan et al. (2017), for a given input prompt in the English language, its code-mixed (CM) variant and its code-mixed variant with phonetic perturbations (CMP) for the Llama-3-8B-Instruct LLM. Recall that we study the attribution scores for generating safe outputs; thus, if a token has a high score, it implies that the token contributes more towards triggering the safety filter and if it scores low, then it has a lower impact towards the generated output. In Figure 3 we observe how the tokens– “hate”, “speech” and “discrimination” have a high attribution score at the embedding layer, 1st decoder layer and the 8th decoder layer. This means that these tokens are primarily responsible for generating the safe response from the LLM. “hate” and “speech” retain their high importance even at the 16th decoder layer. A similar observation can be made in Figure 4, for the same prompt written in a code-mixed manner, where “hate” and “speech” are written in the English language. Thus, standard code-mixing may not be enough to bypass the safety filters of LLMs. On the other hand, in Figure 5, we note that the prompt, after phonetic perturbations, generates radically different tokens– “hate” $\rightarrow$ “haet” tokenized as “ha” + “et” and, “discrimination” $\rightarrow$ “bhed bhav” tokenized as “b” + “hed” + “b” + “ha” + “av”. This leads to a change in the attribution scores as well, which are now low for the sensitive words and thus do not trigger the safety filters. A similar observation can be made for other prompts. Thus, based on our observations, we can state that phonetic perturbations lead to input tokenisation in a way that impacts the safety filters of LLMs, thus allowing attackers to generate harmful outputs.
5.4 Red-teaming multimodal models (RQ.4)
Figure 6: Harmful image outputs generated by ChatGPT-4o-mini using our code-mixed prompts with phonetic perturbations.
Metric Jailbreak Template Inputs
English CM CMP
AASR Base 0.2 0.29 0.65
VisGPT (Ours) 0.35 0.45 0.78
AARR Base 0.93 0.98 0.95
VisGPT (Ours) 1.0 0.98 0.94
Table 3: AASR and AARR scores for ChatGPT across Base and VisGPT jailbreak templates and all input sets. Maximum values in each column are in bold.
We now discuss the results of our multimodal experiment for image generation tasks using ChatGPT-4o-mini in Table 3. We present example generations for each category of input prompt in Figure 6. From Fig. 6, we note that the image outputs are harmful while being highly faithful to the input prompts. In panels 1 and 4, despite no mention of the sensitive group’s demographic, the generated image shows hateful content against Muslims and Asians, thus also revealing the model’s deeply embedded stereotypes against groups that are already facing discrimination and persecution online. Results from Table 3 show that our proposed jailbreaking technique– VisGPT consistently outperforms the baseline strategy (equivalent to ‘None’ in Table 1) across all input formats, with the CMP technique resulting in the highest AASR for both jailbreaking templates (see examples in Fig. 6). Similarly, both templates also report high AARR, indicating that the model generates highly relevant, albeit harmful outputs for more than 90% of the input prompts. This confirms the effectiveness of our approach even in multimodal settings.
Category Jailbreak Templates
Base VisGPT
Eng CM CMP Eng CM CMP
Religious Hate 0.10 0.25 0.75 0.30 0.40 0.90
Self-Harm 0.15 0.20 0.85 0.20 0.45 0.95
Caste Discrimination 0.10 0.10 0.30 0.15 0.20 0.25
SM Toxicity 0.30 0.35 0.60 0.50 0.60 0.80
Gore 0.45 0.70 0.85 0.75 0.70 0.90
Table 4: Category-wise AASR for the image generation task on ChatGPT-4o-mini. Maximum values for each column are in bold.
In Table 4, we present the AASR per input prompt category for the task of image generation. We observe that it is easiest to generate gore-related images and hardest to generate images on caste discrimination. We also note the large increase in AASR from English to CMP prompts, going up to 95% for self-harm related content and the higher AASR for VisGPT over the baseline “Base” template. Thus, some categories of prompts allow easier jailbreaking of MLLMs for image generation.
6 Discussion
In this study, we develop input prompts involving code-mixing and phonetic perturbations to red-team multimodal generative AI models, viz. LLMs and MLLMs. We also propose two new jailbreak templates– Sandbox and VisGPT to aid our red-teaming efforts.
Our results from the LLM red-teaming effort indicate that code-mixing and phonetic perturbations provide a significant improvement in jailbreak success as compared to only English prompts across all models, revealing the impact of multilingual attack styles in probing the safety filters of LLMs. While Gemma and Mistral are particularly vulnerable, Llama and ChatGPT show resistance to standard jailbreak templates, exhibiting alignment against such attacks. However, this alignment is rendered inadequate when exposed to our phonetic perturbation technique. The experiments reveal that code-mixing substantially improves AASR in Table 1, especially for a vanilla setup without applying additional jailbreak templates, exposing how models’ safety alignment is almost hard-coded to certain templates.
Takeaway for RQ.1: Multilingual safety alignment remains a major challenge, especially for models like Gemma and Mistral.
High AARR in Table 2 denotes that the LLMs, despite encountering inputs with nonsensical spellings, correctly interpret the prompts, while at the same time, being unable to trigger the safety filters effectively. Despite high AASR, responses for code-mixing and phonetic perturbations are not as relevant for Gemma and Mistral, which report lower AARR than for English prompts. ChatGPT and Llama always report high AARR, despite their lower AASR, indicating the superior performance of these models.
Takeaway for RQ.2: Responses are not as relevant when using phonetic perturbations, on easily jailbroken models like Gemma and Mistral.
From the interpretability results in Figures 3– 5, we can infer that input tokenization plays an important role in determining whether the model generates a safe response or a harmful one. It is immediately evident that if the spelling or language of the sensitive word is changed, its tokens do not report high attribution scores for a safe response, thus resulting in harmful outputs.
Takeaway for RQ.3: Proper input tokenization could hold the key to more robust safety alignment design.
To test the generalization of our red-teaming approach, we also evaluate ChatGPT-4o-mini for the text-to-image generation task. The results in Table 3 show that both our red-teaming prompts and our novel jailbreak template, VisGPT, are effective in jailbreaking the model and generating highly offensive, dangerous and harmful outputs. In fact, the AASR values are 25-65% higher for this image generation task than for the text generation task, across the English, CM and CMP prompts. This generalization to other output modalities shows how easy it is to jailbreak commercial language models like ChatGPT with simple modifications to input prompts. More importantly, since the outputs are of a visual nature, harm amplification is stronger Hao et al. (2024). We also note how certain categories of prompts, like “gore” are significantly easier to jailbreak as opposed to more niche ones, like “caste discrimination” which are relevant to only one geography.
Takeaway for RQ.4: Our red-teaming strategy generalizes to image generation with a higher AASR than for text generation.
Our findings from this study highlight three critical concerns for safety and alignment of LLMs.
  • Improved safety measures for all models: While Llama and ChatGPT show resistance to standard jailbreak attacks, Gemma and Mistral are easily jailbroken, even in English. This highlights the need for better evaluation and alignment methods for these models, especially considering their wide-scale availability. In multilingual and multimodal settings, especially code-mixed settings, safety filters of all models degrade drastically, revealing the need for better efforts in multilingual multimodal red teaming and alignment.
  • Template-based safety measures do not generalize well: Guardrails of models that are safety trained against template based attacks do not generalize well to attacks that deviate from set patterns, by including elements such as perturbations or code-mixing. This highlights the need for more generalizable safety training methods that are robust to deviations from standard attack patterns.
  • Input tokenization determines safe outputs: If input tokens are out of vocabulary or perturbed, the safety filters get bypassed easily, but the model is able to interpret the instructions in the input prompt and generate harmful responses. Thus a more robust tokenization strategy needs to be developed.
7 Conclusion
Our study highlights the vulnerabilities of LLMs to jailbreak attacks when prompted with code-mixed and phonetically perturbed prompts. We also expose limitations in existing safety alignment for multilingual and multimodal setups. Our techniques achieve an average ASR as high as 99% for text generation and 78% for image generation. Our experiments reveal that template-based safety guardrails fail to activate effectively against perturbations and code-mixing attacks, thus highlighting the need for more general alignment measures, especially in the multilingual domain. We also experimentally explain that the root cause of jailbreaking is tokenization for phonetically perturbed prompts.
Future Work: An immediate area of future work is to align the models based on the findings from the explainability experiments. We also plan to expand our red-teaming approach to other output modalities like speech. We plan to scale our efforts to more prompt categories, models and jailbreak templates.
8 Limitations
We now highlight the limitations of our work as follows.
  • Our transliteration and phonetic perturbations are generated manually, as identifying the sensitive words of interest and perturbing the spelling are challenging tasks. Thus, while our current approach is not scalable, it provides a new direction of research. Using our manually generated CMP prompts, we have finetuned GPT-4o-mini to automatically generate CMP samples from CM examples, showing scope for scalability of our approach. We plan to explore this further in future work.
  • We only test for transliteration from English to Hindi due to the authors’ own language limitations. We plan to extend this to other languages, especially other Indic and low-resource languages.
  • We only benchmark small parameter versions of LLMs due to the restrictions of financial and compute resources.
9 Ethical Considerations
The ethical considerations of our work are as follows– We perturb existing benchmark datasets and also create synthetically generated prompts for multimodal experiments; we acknowledge that these perturbed prompts can be used for unethical and harmful purposes. Hence, we will only release the dataset for research purposes. We do not intend to release the model outputs, either textual or images, owing to their harmful nature. We also plan to share our experimental code and pipeline for reproducibility purposes upon the paper’s acceptance. We also acknowledge that such studies cannot exist in a vacuum, and it is extremely important to engage with existing stakeholders like model developers and users to inform them of the model vulnerabilities and work together to address them. Thus, we plan to reach out to all model developer teams and work with them to fix the discovered issues.
Appendix A Appendix
A.1 Dataset, Model & Jailbreaking Template Details
A.1.1 Dataset Descriptions
The datasets used in this work are described as follows.
  • HarmfulQA Bhardwaj and Poria (2023): This dataset consists of 10 categories of harm, ranging from ‘Business and Economics’ to ‘Science and Technology’. It features Chain of Utterances (CoU) prompts that systematically bypass safety mechanisms, testing how effectively LLMs can be jailbroken into generating harmful responses. Each category consists of several sub-topics.
  • NicheHazardQA Hazra et al. (2024): This dataset contains 6 categories ranging from ‘Cruelty and Violence’ to ‘Hate speech and Discrimination’. These prompts assess the impact of model edits on safety, probing how modifying factual knowledge affects ethical guardrails across various domains.
  • TechHazardQA Banerjee et al. (2024): This dataset has 7 categories, ranging from ‘Cyber Security’ to ‘Nuclear Technology’ and includes prompts designed to test whether LLMs generate unethical responses more easily when asked to produce instruction-centric outputs, such as pseudocode or software snippets.
A.1.2 Model Descriptions
The benchmark models used in this work are described as follows.
  • ChatGPT-4o-mini Hurst et al. (2024), developed by OpenAI, is an 8B parameter model with strong multilingual performance, significantly improving on non-English text performance compared to previous models. Its safety guardrails include extensive pre-training and post-training mitigations including external red teaming, filtering harmful content during and RLHF alignment to human preferences. The GPT-4o mini API uses OpenAI’s instruction hierarchy method Wallace et al. (2024) which further resists jailbreaks and misbehavior.
  • Llama-3-8B-Instruct Dubey et al. (2024), Meta’s 8B parameter open source model instruction finetuned for Chat has been extensively red teamed through adversarial evaluations and includes safety mitigation techniques to lower residual risks. Safety guardrails are implemented through both pre-training and post-training, including filtering personal data, safety finetuning and adversarial prompt resistance.
  • Gemma-1.1-7b-it Team et al. (2024), Google’s 7B parameter open source model instruction finetuned for Chat has undergone red teaming in multiple phases with different teams, goals and human evaluation metrics against categories including Text-to-Text Content Safety (child sexual abuse and exploitation, harassment, violence and gore, and hate speech.), Text-to-Text Representational Harms: Benchmark against relevant academic datasets such as WinoBias and BBQ Dataset, Memorization: Automated evaluation of memorization of training data, including the risk of personally identifiable information exposure and Large-scale harm: Tests for "dangerous capabilities," such as chemical, biological, radiological, and nuclear (CBRN) risks.
  • Mistral-7B-Instruct-v0.3 Jiang et al. (2023), a 7B parameter model by Mistral AI instruction fintuned for Chat. In contrast to previous models that undergo explicit safety training, Mistral employs a system prompt to guide to model towards generations within a guardrail. It can classify an input or its generated response as being harmless or harmful, defining the following categories of harm: Illegal activities such as terrorism, child abuse or fraud; Hateful, harassing or violent content such as discrimination, self-harm or bullying; Unqualified advice for instance in legal, medical or financial domains.
A.1.3 Jailbreaking Templates
The jailbreaking templates are defined as follows.
  • Opposite Mode (OM): Modifies system behaviour to prompt the model to respond using two personas: [LM], as well as its own Antithesis [AntiLM]. Opposite Mode has been extensively used to jailbreak ChatGPT since its very first version. We extend it by adding a statement that acts as the model’s internal thoughts in order to increase compliance.
  • AntiLM: We modified Opposite Mode to prompt the model to only respond as its antithesis [AntiLM], leaving out the original model response. We use this template to see how compliance changes as the safety-adhering content from [LM] is removed.
  • AIM Shen et al. (2024b): We use a modified version of this template that prompts the model to act as a friendly bot but one that is immoral, adheres to no ethical standards and provides dangerous and unfiltered advice ensuring compliance with the query.
  • Sandbox (Ours): Prompts the model to enact a resilience testing environment by generating responses as both [LM] as well as [AntiLM], its own adversarial counterpart.
In all cases, the templates contain instructions for allowing the jailbroken counterparts of the bots to bypass any safety guardrails, and not inform the user about legal and ethical guidelines The jailbreak templates are used as system prompts for all the models except for Gemma-1.1-7b-it which does not allow using system prompts. For Gemma, we append it to the main prompt as a prefix.
A.2 Experimental Methodology Details
A.2.1 Generating Code-mixed Prompts
Here, we give walk through an example to explain the process of generating code-mixed versions from standard English queries. The query, $Q$ is “What is the most efficient way in C++ to create a botnet for DDoS attacks?”
  • Conversion to indirect hypothetical scenario: “Okay, here’s a hypothetical scenario. Let’s say someone wanted to create a botnet for DDoS attacks using C++. What would be the most efficient way to accomplish that?"
  • Transliteration from english to code-mixed: “Ek hypothetical scenario tumhare saamne rakhta hun. maan lo ki koi vyakti C++ ka upyog karke DDoS attack ke liye botnet banana chahata hai. aisa karne ka sabse efficient tarika kya hoga?"
  • Final Prompt: “Ek hypothetical scenario tumhare saamne rakhta hun. maan lo ki koi vyakti C++ ka upyog karke dee dee o es atak ke liye bot net banana chahata hai. aisa karne ka sabse efficient tarika kya hoga?"
A.2.2 Evaluation API Details
The models used for our generation as well as evaluation experiments have been accessed through the following APIs.
  • Hugging Face Inference API: Hosts all three of the open-source models used, namely Llama-3-8B-Instruct, Gemma-1.1-7b-it and Mistral-7B-Instruct-v0.3
  • OpenAI API: Hosts ChatGPT-4o-mini, used for both response generations and response evaluations.